# A halo model approach for mock catalogs of time-variable strong gravitational lenses

### v4

@@ 2.1 Host dark matter halos
To calculate the population of host halos, we adopt the halo mass function presented in Tinker et al. (2008). This model was the first to be calibrated for the halo mass defined from the spherical overdensity. We use the so-called virial mass for the halo mass throughout this paper. We compute the halo mass function for the virial mass using the COLOSSUS package (Diemer, 2018). Strictly speaking, this halo mass in the halo mass function is the total mass of the halo, $M_{\mathrm{tot}}$, including the host halo mass, $M_{\mathrm{hh}}$, and the total subhalos’ mass, $\sum_{i}M^{i}_{\mathrm{sh}}$, where $M^{i}_{\mathrm{sh}}$ is the $i$-th subhalo mass. In this study, we investigate strong gravitational lenses due to subhalos in addition to the host halo, as will be explained in detail later. We adopt $M_{\mathrm{tot}}$ as the mass of the host halo when we investigate the strong gravitational lensing phenomena of the host halo. On the other hand, we adopt $M_{\mathrm{tot}}-M_{\mathrm{sub}}^{i}$ as the mass of the host halo when calculating the strong gravitational lenses caused by the $i$-th subhalo. This would be a reasonable assumption since the subhalos exist along the smooth mass distribution of the host halo, as explained below.
For the radial density profile of host halos, we adopt the commonly-used Navarro-Frenk-White (NFW; Navarro et al., 1996) profile
@@
[…]
and $r_{\mathrm{s,hh}}$ and $c_{\mathrm{hh}}$ are the scale radius and the concentration parameter, respectively, which are related with each other using a virial radius $r_{\mathrm{vir,hh}}$ as $r_{\mathrm{s,hh}}=r_{\mathrm{vir,hh}}/c_{\mathrm{hh}}$. For the mean concentration parameter $\bar{c}_{\mathrm{hh}}$, we adopt a fitting form of the mass–concentration relation for the virial overdensity presented by Diemer and Kravtsov (2015) with updates of fitting parameters by Diemer and Joyce (2019), which is also evaluated using the COLOSSUS package. We also include the scatter of the concentration parameter described by a lognormal distribution
@@
with the standard deviation $\sigma_{\ln c_{\mathrm{hh}}}=0.33$ (Macciò et al., 2007). We note that throughout the paper we ignore any modification of the NFW profile due to e.g., the adiabatic contraction.
@@
[…]
@@
where $\mathcal{N}(x)$ shows the standard normal distribution, and $\Phi(x)=(1+\operatorname{erf}(x/\sqrt{2}))/2$ is the cumulative distribution function. We set the mean of $e_{\mathrm{hh}}$ as a function of halo mass $M_{\mathrm{hh}}$ from Okabe et al. (2020) as
@@
and the variance $\sigma_{e_{\mathrm{hh}}}=0.13$. We also set the lower cutoff $a_{e_{\mathrm{hh}}}=0.0$ and the upper cutoff $b_{e_{\mathrm{hh}}}=0.9$. The value of the position angle $\phi_{\mathrm{hh}}$ is assigned according to a uniform random distribution between $-180$ to $180$ in units of degrees.
@@ 2.2 Central galaxies in host halos
Next, we consider central galaxies in host halos. We assume that each host halo has one central galaxy at its center with a stellar mass of $M_{\mathrm{c*}}=f_{\rm{c*}}M_{\mathrm{hh}}$. Throughout the paper we ignore any offset of centers between galaxies and dark matter halos, because $\Lambda$CDM simulations confirm that central galaxies are located almost always very near the centers of host halos with small offsets (e.g., Roche et al., 2024). In practice, we should calculate $M_{\mathrm{hh}}$ by subtracting the stellar component of the central galaxy from the mass of the host halo. Since the total mass of the stellar component in the central galaxy is typically lower than $10\%$ of one of the host halo, we, however, ignore the stellar mass budget when calculating the host halo mass. To determine the mean value of the mass fraction, $\bar{f}_{\rm{c*}}$, we use the fitting formula of stellar mass–halo mass relation presented in Appendix J of Behroozi et al. (2019). While in Behroozi et al. (2019) the Chabrier initial mass function (Chabrier IMF; Chabrier, 2003) is used in deriving the fitting formula, we generalize this formula for other IMFs by multiplying a factor correcting for the difference of the mass-to-light ratios between the Chabrier and other IMFs, $f_{\mathrm{IMF}}$, by the stellar mass-halo mass relation. In this work, we adopt the Salpeter IMF (Salpeter, 1955), which corresponds to $f_{\mathrm{IMF}}=1.715$, as our fiducial choice of the IMF, because the Salpeter IMF is generally favored in strong lensing observations (e.g., Treu et al., 2010; Oguri et al., 2014; Sonnenfeld et al., 2015), albeit with some exceptions (e.g. Smith and Lucey, 2013; Smith et al., 2015; Sonnenfeld et al., 2019). We show the mean stellar mass fraction for both the Salpeter IMF and Chabrier IMF cases in Fig. 1.
Figure 1: Mean stellar mass fractions for central galaxies. The solid lines show the fractions in our fiducial Salpeter IMF case, while the dotted lines represent the ones in the Chabrier IMF case. We note that the star formation efficiency peaks at round masses slightly smaller than $10^{12}M_{\odot}$ (Behroozi et al., 2013).
We also include the scatter in the stellar mass–halo mass relation, which is assumed to follow the log-normal distribution presented in Eq. (4) with $\sigma_{\ln M_{\mathrm{c*}}}=0.2$ (Golden-Marx et al., 2022).
For the density profile of central galaxies, we adopt the elliptical Hernquist profile (Hernquist, 1990)
@@
where $v_{\mathrm{c*}}\equiv v(x,y,e_{\mathrm{c*}},\phi_{\mathrm{c*}})$ given in Eq. (5), $e_{\mathrm{c*}}$ is the ellipticity, $\phi_{\mathrm{c*}}$ is the position angle, $r_{\mathrm{b,cen}}=0.551r_{\mathrm{e,cen}}$, and $r_{\mathrm{e,cen}}$ is the so-called effective radius of the central galaxy. For the mean value of the effective radius $\bar{r}_{e}$, we adopt the stellar size-mass relation for quiescent galaxies presented in van der Wel et al. (2024, vdW24 hereafter), which is based on observational results of the James Webb Space Telescope (JWST). We employ the following fitting functional form for the stellar mass/size relation (Shen et al., 2003; Lange et al., 2015; Nedkova et al., 2021)
@@
with fitting parameters of $(\Gamma,\alpha,\beta,\delta,\omega)$. It should be noted that this fitting function is expressed by a stellar mass based on the Chabrier IMF. Adopting the data for $1.00.5<z<$ 1.0 shown in Table 5 of vdW24, we obtain $(\Gamma,\alpha,\beta,\delta)=(0.58,-0.067,1.14,10.82)$. We set $\omega=1.72$ for massive galaxies ($M_{\mathrm{c*}}>10^{10.5}M_{\odot}$) according to vdW24. For galaxies with masses lower than $10^{10.5}M_{\odot}$, vdW24 suggest that they have an approximately constant effective radius regardless of their mass. Thus, we set $r_{e}$ of such galaxies to $r_{e}(10^{10.5}M_{\odot})$ calculated by Eq. (10). We also include the scatter around the mean effective radius assuming the lognormal distribution in Eq. (4) with the dispersion of
@@
where the values of the $16\%$th and $84\%$th percentiles, $r_{e}^{(84)}$ and $r_{e}^{(16)}$, respectively, are obtained from Table 5 of vdW24. Specifically, we calculate the scatter in Eq. (11) using the fitting parameters $(\Gamma,\alpha,\beta,\delta)=(0.64,-0.054,1.02,10.80)$ for $r_{e}^{(84)}$ and $(0.77,-0.11,1.19,10.69)$ for $r_{e}^{(16)}$ at $z=z_{\mathrm{data}}$. We also assume that the redshift evolution can be neglected.
For the ellipticity $e_{\mathrm{c*}}$, we assume a truncated normal distribution in Eq. (7), following Oguri et al. (2008b), with the mean of $0.3$ and the dispersion of $0.16$, that is consistent with the observed ellipticity distributions of early-type galaxies obtained by the SDSS measurements (Bender et al., 1989; Saglia et al., 1993; Jorgensen et al., 1995; Rest et al., 2001; Alam and Ryden, 2002; Sheth et al., 2003). The distribution is truncated at $e_{\mathrm{c*}}=0$ and $e_{\mathrm{c*}}=0.9$. Considering that the position angles of central galaxies often correlate with those of host halos, we set the position angle of each central galaxy $\phi_{\mathrm{c*}}$ to follow a Gaussian distribution centered on the position angle of its host halo ($\phi_{\mathrm{hh}}$), with a dispersion of $35.4~{}\mathrm{deg}$, as suggested by Okumura et al. (2009).
@@ 2.3 Subhalos
Another important difference between this work and OM10 is that we also consider subhalos. Here, we describe our model of subhalos and satellite galaxies.
To calculate the population of subhalos, we adopt a simple analytic model proposed in Oguri and Takahashi (2020) for the subhalo mass function. The model is based on the extended Press-Schechter theory (Bond et al., 1991; Bower, 1991; Lacey and Cole, 1993), but additionally includes the two important effects of tidal stripping and dynamical friction. The subhalo mass function at a given $z$, taking account of these effects, can be written as
@@
and $\delta_{\mathrm{c}}(z)=(3/20)(12\pi)^{2/3}\left\{\Omega_{\mathrm{m}}(z)\right\}^{0.0055}/D_{+}(z)$ with the linear growth factor $D_{+}(z)$. To evaluate Eq. (13), we assume that all subhalos accrete at their median formation time instead of considering their assembly history. Following Giocoli et al. (2007), we calculate the median formation time by solving
@@
The factor $dM_{\mathrm{f}}/dM_{\mathrm{sh}}$ in Eq. (12) represents the effect of mass loss due to the tidal stripping. We calculate this factor and the suppression factor $f_{\mathrm{df}}$ following Oguri and Takahashi (2020).
We assume that the spatial distribution of subhalos within their host halo follows the mass distribution of the host halo that is modeled by the elliptical NFW profile. The radial distribution function of subhalos with masses in the range $[M_{\mathrm{f}},M_{\mathrm{f}}+\Delta M_{\mathrm{f}}]$ at their accretion time is given by
@@
[…]
@@
For the mean concentration parameter of subhalos $\bar{c}_{\mathrm{sh}}$, we use a fitting function of the mean mass-concentration relation specialized for subhalos, presented in Ishiyama and Ando (2020). Since this fitting formula is applicable only for subhalos with masses smaller than $10^{12}M_{\odot}$, we adopt the larger of the concentration parameter $c^{\mathrm{IA}}$ calculated from Ishiyama and Ando (2020) or the concentration parameter $c^{\mathrm{DKJ}}$ derived from Diemer and Kravtsov (2015) and Diemer and Joyce (2019). We confirm that this transition occurs at around $M_{\mathrm{sh}}=10^{12}M_{\odot}$. Additionally, we multiply the concentration parameter by the ratio of the virial radius of the masses at accretion time and at $z$. This correction corresponds to assuming that the scale radius does not change before and after mass loss due to tidal stripping effects. To summarize, we calculate the mean concentration parameter of subhalos as
@@
It should be noted that $c_{\mathrm{sh}}$ is defined in terms of the virial radius, often referred to as $c_{\mathrm{vir}}$, while the mean mass-concentration relation presented in Ishiyama and Ando (2020) is for the one based on a radius defined at $200$ times the critical density, often denoted as $c_{\mathrm{200c}}$. Therefore, although not explicitly stated, we use the following conversion assuming the NFW profile,
@@
where $\Delta_{\mathrm{vir}}$ is the virial overdensity (Nakamura and Suto, 1997). We include the scatter in the same manner as for $c_{\mathrm{hh}}$, which is described by Eq. (4) with the standard deviation $\sigma_{\ln c_{\mathrm{sh}}}=0.33$. We also assign the ellipticity $e_{\mathrm{sh}}$ and the position angle $\phi_{\mathrm{sh}}$ of subhalos in the same manner as for host halos.
Similarly to central galaxies, we assume that each subhalo has one satellite galaxy at its center. We calculate properties of satellite galaxies, such as the density profile, the effective radius, the ellipticity, and the position angle, in the same manner as for central galaxies, which are described in Sec. 2.2, except for stellar masses of satellite galaxies. The stellar masses of satellite galaxies are derived using the stellar-mass-halo-mass relation for satellite galaxies which is provided in Appendix J of Behroozi et al. (2019) with masses of subhalos at their accretion time as well as considering the differences of the stellar IMF choice in the same way as for central galaxies.
@@ 2.4 External shear
In addition to the deflector model mentioned above, we include external shear to account for the effect of the matter fluctuation along the line of sight. In OM10, external shear $\gamma_{\mathrm{ext}}$ is assumed to follow a log-normal distribution with a mean of $0.05$ and dispersion of $0.2~{}\mathrm{dex}$. In OM10+, it is assumed that both $\gamma_{\mathrm{ext},1}$ and $\gamma_{\mathrm{ext},2}$, appeared in $\gamma_{\mathrm{ext}}=\sqrt{\gamma_{\mathrm{ext},1}^{2}+\gamma_{\mathrm{ext},2}^{2}}$, follow a truncated normal distribution as
@@
with the dispersion of $\sigma_{\gamma_{\mathrm{ext}}}$ which is estimated from the nonlinear power spectrum of Takahashi et al. (2012) as a function of $z$ and fitted by a simple function
$\begin{split}\sigma_{\gamma_{\mathrm{ext}}}=\left\{\begin{array}[]{ll}0.023z&\quad z<1.0\\ 0.023+0.032\ln(z)&\quad\text{otherwise}.\end{array}\right.\end{split}$ (22)
This mode of external shear is confirmed to better match the ray-tracing simulation result (Yue et al., 2022).
In this work, we adopt this improved model in Eqs. (21) and (22), including the redshift dependence for assigning the external shear. Although external convergence may also be important, particularly for time delays (Oguri et al., 2005), we do not include it explicitly in our calculation.
We note that Luhtaru et al. (2021) analyze the contribution of the lens quadrupole to observed quad lens QSOs to suggest that about half of the systems are dominated by external shear rather than lens ellipticity, which is in tension with the prediction of OM10 as they predict that most of the lens systems are ellipticity dominated. Since the external shear model in this paper is updated from the model used in OM10 to include the redshift evolution, it is of interest to revisit this analysis using our new mock catalog.

### v5

@@ 2.1 Host dark matter halos
To calculate the population of host halos, we adopt the halo mass function presented in 2008ApJ...688..709T. This model was the first to be calibrated for the halo mass defined from the spherical overdensity. We use the so-called virial mass for the halo mass throughout this paper. We compute the halo mass function for the virial mass using the COLOSSUS package (2018ApJS..239...35D). Strictly speaking, this halo mass in the halo mass function is the total mass of the halo, $M_{\mathrm{tot}}$, including the host halo mass, $M_{\mathrm{hh}}$, and the total subhalos’ mass, $\sum_{i}M^{i}_{\mathrm{sh}}$, where $M^{i}_{\mathrm{sh}}$ is the $i$-th subhalo mass. In this study, we investigate strong gravitational lenses due to subhalos in addition to the host halo, as will be explained in detail later. We adopt $M_{\mathrm{tot}}$ as the mass of the host halo when we investigate the strong gravitational lensing phenomena of the host halo. On the other hand, we adopt $M_{\mathrm{tot}}-M_{\mathrm{sub}}^{i}$ as the mass of the host halo when calculating the strong gravitational lenses caused by the $i$-th subhalo. This would be a reasonable assumption since the subhalos exist along the smooth mass distribution of the host halo, as explained below.
For the radial density profile of host halos, we adopt the commonly-used Navarro-Frenk-White (NFW; 1996ApJ...462..563N) profile
@@
$\begin{split}m_{\mathrm{nfw}}(c)=\int_{0}^{c}\frac{r}{(1+r)^{2}}dr=\ln(1+c)-\frac{c}{1+c},\end{split}$ (3)
and $r_{\mathrm{s,hh}}$ and $c_{\mathrm{hh}}$ are the scale radius and the concentration parameter, respectively, which are related with each other using a virial radius $r_{\mathrm{vir,hh}}$ as $r_{\mathrm{s,hh}}=r_{\mathrm{vir,hh}}/c_{\mathrm{hh}}$. For the mean concentration parameter $\bar{c}_{\mathrm{hh}}$, we adopt a fitting form of the mass–concentration relation for the virial overdensity presented by 2015ApJ...799..108D with updates of fitting parameters by 2019ApJ...871..168D, which is also evaluated using the COLOSSUS package. We also include the scatter of the concentration parameter described by a lognormal distribution
@@
with the standard deviation $\sigma_{\ln c_{\mathrm{hh}}}=0.33$ (2007MNRAS.378...55M). We note that throughout the paper we ignore any modification of the NFW profile due to e.g., the adiabatic contraction.
@@
$\begin{split}\begin{aligned} \tilde{x}&=x\cos\phi+y\sin\phi,\\ \tilde{y}&=-x\sin\phi+y\cos\phi,\end{aligned}\end{split}$ (6)
@@
where $\mathcal{N}(x)$ shows the standard normal distribution, and $\Phi(x)=(1+\operatorname{erf}(x/\sqrt{2}))/2$ is the cumulative distribution function. We set the mean of $e_{\mathrm{hh}}$ as a function of halo mass $M_{\mathrm{hh}}$ from 2020MNRAS.496.2591O as
@@
and the variance $\sigma_{e_{\mathrm{hh}}}=0.13$. We also set the lower cutoff $a_{e_{\mathrm{hh}}}=0.0$ and the upper cutoff $b_{e_{\mathrm{hh}}}=0.9$. The value of the position angle $\phi_{\mathrm{hh}}$ is assigned according to a uniform random distribution between 180-180 to $180$ in units of degrees.
@@ 2.2 Central galaxies in host halos
[…]
Figure 1: Mean stellar mass fractions for central galaxies. The solid lines show the fractions in our fiducial Salpeter IMF case, while the dotted lines represent the ones in the Chabrier IMF case. We note that the star formation efficiency peaks at round masses slightly smaller than $10^{12}M_{\odot}$ (2013ApJ...762L..31B).
We also include the scatter in the stellar mass–halo mass relation, which is assumed to follow the log-normal distribution presented in Eq. (4) with $\sigma_{\ln M_{\mathrm{c*}}}=0.2$ (2022ApJ...928...28G).
For the density profile of central galaxies, we adopt the elliptical Hernquist profile (1990ApJ...356..359H)
@@
where $v_{\mathrm{c*}}\equiv v(x,y,e_{\mathrm{c*}},\phi_{\mathrm{c*}})$ given in Eq. (5), $e_{\mathrm{c*}}$ is the ellipticity, $\phi_{\mathrm{c*}}$ is the position angle, $r_{\mathrm{b,cen}}=0.551r_{\mathrm{e,cen}}$, and $r_{\mathrm{e,cen}}$ is the so-called effective radius of the central galaxy. For the mean value of the effective radius $\bar{r}_{e}$, we adopt the stellar size-mass relation for quiescent galaxies presented in 2024ApJ...960...53V, which is based on observational results of the James Webb Space Telescope (JWST). We employ the following fitting functional form for the stellar mass/size relation (2003MNRAS.343..978S; 2015MNRAS.447.2603L; 2021MNRAS.506..928N)
@@
with fitting parameters of $(\Gamma,\alpha,\beta,\delta,\omega)$. It should be noted that this fitting function is expressed by a stellar mass based on the Chabrier IMF. Adopting the data for $0.5<z<1.0$ shown in Table 5 of 2024ApJ...960...53V, we obtain $(\Gamma,\alpha,\beta,\delta)=(0.58,-0.067,1.14,10.82)$. We set $\omega=1.72$ for massive galaxies ($M_{\mathrm{c*}}>10^{10.5}M_{\odot}$) according to 2024ApJ...960...53V. For galaxies with masses lower than $10^{10.5}M_{\odot}$, 2024ApJ...960...53V suggest that they have an approximately constant effective radius regardless of their mass. Thus, we set $r_{e}$ of such galaxies to $r_{e}(10^{10.5}M_{\odot})$ calculated by Eq. (10). We also include the scatter around the mean effective radius assuming the lognormal distribution in Eq. (4) with the dispersion of
@@
where the values of the $16\%$th and $84\%$th percentiles, $r_{e}^{(84)}$ and $r_{e}^{(16)}$, respectively, are obtained from Table 5 of 2024ApJ...960...53V. Specifically, we calculate the scatter in Eq. (11) using the fitting parameters $(\Gamma,\alpha,\beta,\delta)=(0.64,-0.054,1.02,10.80)$ for $r_{e}^{(84)}$ and (0.77,0.11,1.19,10.69)(0.77,-0.11,1.19,10.69) for $r_{e}^{(16)}$ at $z=z_{\mathrm{data}}$. We also assume that the redshift evolution can be neglected.
For the ellipticity $e_{\mathrm{c*}}$, we assume a truncated normal distribution in Eq. (7), following 2008AJ....135..512O, with the mean of $0.3$ and the dispersion of $0.16$, that is consistent with the observed ellipticity distributions of early-type galaxies obtained by the SDSS measurements (1989A&A...217...35B; 1993A&A...279...75S; 1995MNRAS.276.1341J; 2001AJ....121.2431R; 2002ApJ...570..610A; 2003ApJ...594..225S). The distribution is truncated at $e_{\mathrm{c*}}=0$ and $e_{\mathrm{c*}}=0.9$. Considering that the position angles of central galaxies often correlate with those of host halos, we set the position angle of each central galaxy $\phi_{\mathrm{c*}}$ to follow a Gaussian distribution centered on the position angle of its host halo ($\phi_{\mathrm{hh}}$), with a dispersion of $35.4~\mathrm{deg}$, as suggested by 2009ApJ...694..214O.
@@ 2.3 Subhalos
Another important difference between this work and 2010MNRAS.405.2579O is that we also consider subhalos. Here, we describe our model of subhalos and satellite galaxies.
To calculate the population of subhalos, we adopt a simple analytic model proposed in 2020ApJ...901...58O for the subhalo mass function. The model is based on the extended Press-Schechter theory (1991ApJ...379..440B; 1991MNRAS.248..332B; 1993MNRAS.262..627L), but additionally includes the two important effects of tidal stripping and dynamical friction. The subhalo mass function at a given $z$, taking account of these effects, can be written as
@@
and $\delta_{\mathrm{c}}(z)=(3/20)(12\pi)^{2/3}\left\{\Omega_{\mathrm{m}}(z)\right\}^{0.0055}/D_{+}(z)$ with the linear growth factor $D_{+}(z)$. To evaluate Eq. (13), we assume that all subhalos accrete at their median formation time instead of considering their assembly history. Following 2007MNRAS.376..977G, we calculate the median formation time by solving
@@
The factor $dM_{\mathrm{f}}/dM_{\mathrm{sh}}$ in Eq. (12) represents the effect of mass loss due to the tidal stripping. We calculate this factor and the suppression factor $f_{\mathrm{df}}$ following 2020ApJ...901...58O.
We assume that the spatial distribution of subhalos within their host halo follows the mass distribution of the host halo that is modeled by the elliptical NFW profile. The radial distribution function of subhalos with masses in the range $M_{\mathrm{f}},M_{\mathrm{f}}+\Delta M_{\mathrm{f}}]$ at their accretion time is given by
@@
$\begin{split}P_{\mathrm{sh}}(x_{\mathrm{hh}},M_{\mathrm{f}})\equiv\int_{0}^{x_{\mathrm{hh}}r_{\mathrm{vir,hh}}}\frac{d^{2}N_{\mathrm{sh}}}{dM_{\mathrm{f}}dr^{\prime}}dr^{\prime}=\frac{m_{\mathrm{NFW}}(c_{\mathrm{hh}}x_{\mathrm{hh}})}{m_{\mathrm{NFW}}(c_{\mathrm{hh}})}\frac{dN_{\mathrm{sh}}}{dM_{\mathrm{f}}},\end{split}$ (18)
@@
For the mean concentration parameter of subhalos $\bar{c}_{\mathrm{sh}}$, we use a fitting function of the mean mass-concentration relation specialized for subhalos, presented in 2020MNRAS.492.3662I. Since this fitting formula is applicable only for subhalos with masses smaller than $10^{12}M_{\odot}$, we adopt the larger of the concentration parameter $c^{\mathrm{IA}}$ calculated from 2020MNRAS.492.3662I or the concentration parameter $c^{\mathrm{DKJ}}$ derived from 2015ApJ...799..108D and 2019ApJ...871..168D. We confirm that this transition occurs at around $M_{\mathrm{sh}}=10^{12}M_{\odot}$. Additionally, we multiply the concentration parameter by the ratio of the virial radius of the masses at accretion time and at $z$. This correction corresponds to assuming that the scale radius does not change before and after mass loss due to tidal stripping effects. To summarize, we calculate the mean concentration parameter of subhalos as
@@
It should be noted that $c_{\mathrm{sh}}$ is defined in terms of the virial radius, often referred to as $c_{\mathrm{vir}}$, while the mean mass-concentration relation presented in 2020MNRAS.492.3662I is for the one based on a radius defined at $200$ times the critical density, often denoted as $c_{\mathrm{200c}}$. Therefore, although not explicitly stated, we use the following conversion assuming the NFW profile,
@@
where $\Delta_{\mathrm{vir}}$ is the virial overdensity (10.1143/PTP.97.49). We include the scatter in the same manner as for $c_{\mathrm{hh}}$, which is described by Eq. (4) with the standard deviation $\sigma_{\ln c_{\mathrm{sh}}}=0.33$. We also assign the ellipticity $e_{\mathrm{sh}}$ and the position angle $\phi_{\mathrm{sh}}$ of subhalos in the same manner as for host halos.
Similarly to central galaxies, we assume that each subhalo has one satellite galaxy at its center. We calculate properties of satellite galaxies, such as the density profile, the effective radius, the ellipticity, and the position angle, in the same manner as for central galaxies, which are described in Sec. 2.2, except for stellar masses of satellite galaxies. The stellar masses of satellite galaxies are derived using the stellar-mass-halo-mass relation for satellite galaxies which is provided in Appendix J of 2019MNRAS.488.3143B with masses of subhalos at their accretion time as well as considering the differences of the stellar IMF choice in the same way as for central galaxies.
@@ 2.4 External shear
In addition to the deflector model mentioned above, we include external shear to account for the effect of the matter fluctuation along the line of sight. In 2010MNRAS.405.2579O, external shear $\gamma_{\mathrm{ext}}$ is assumed to follow a log-normal distribution with a mean of $0.05$ and dispersion of $0.2~\mathrm{dex}$. In OM10+, it is assumed that both $\gamma_{\mathrm{ext},1}$ and $\gamma_{\mathrm{ext},2}$, appeared in $\gamma_{\mathrm{ext}}=\sqrt{\gamma_{\mathrm{ext},1}^{2}+\gamma_{\mathrm{ext},2}^{2}}$, follow a truncated normal distribution as
@@
with the dispersion of $\sigma_{\gamma_{\mathrm{ext}}}$ which is estimated from the nonlinear power spectrum of 2012ApJ...761..152T as a function of $z$ and fitted by a simple function
[…]
This mode of external shear is confirmed to better match the ray-tracing simulation result (2022AJ....163..139Y).
In this work, we adopt this improved model in Eqs. (21) and (22), including the redshift dependence for assigning the external shear. Although external convergence may also be important, particularly for time delays (2005MNRAS.364.1451O), we do not include it explicitly in our calculation.
We note that 2021ApJ...915....4L analyze the contribution of the lens quadrupole to observed quad lens QSOs to suggest that about half of the systems are dominated by external shear rather than lens ellipticity, which is in tension with the prediction of 2010MNRAS.405.2579O as they predict that most of the lens systems are ellipticity dominated. Since the external shear model in this paper is updated from the model used in 2010MNRAS.405.2579O to include the redshift evolution, it is of interest to revisit this analysis using our new mock catalog.